# A Similarity Measure for Comparing Conversational Dynamics

### v1

@@ 6.2 New investigation: speakers’ tendencies
The dynamics of a conversation are the results of a joint process involving all speakers. However, similar to how speakers have a tendency to use a specific style across different conversations Welch et al. (2019); Zhang et al. (2020a), they may also have a tendency to engage in certain types of dynamics. In fact, we find evidence to that effect: conversations that share a common speaker have more similar dynamics than those which do not (ConDynS 0.370.370.37 vs. 0.350.350.35, $p<0.001$ according to Mann Whitney U-test, comparing random samples of 2000 pairs of conversations).
@@
Alternatively, prior studies emphasize the critical role of persuasion strategies in debates and their outcomes Braca and Dondio (2023); Orazi et al. (2025). The Challenger, by selecting the strategy, may dictate the dynamics.
To distinguish between these two hypotheses, we design a setup that controls for speaker-related confounds, such as demographics that might otherwise spuriously correlate with both their role in the conversation and their influence (Figure 4). We select speakers who participate in at least four conversations, each started by a different post: two in which they take the role of OP and two in which they are the Challenger. [^11] There are 486486486 such speakers, considering only dyadic conversations that pass our length filters.
We find that the pair in which the speaker is the OP is more similar than the pair in which they are the Challenger (0.410.410.41 vs. $0.37,p<0.01$ according to Wilcoxon signed-rank test). This suggests that conversation dynamics are more likely to follow the tendencies of the (higher-powered) OP than those of the Challenger, supporting the first hypothesis. [^12] We note that this result is not a mere consequence of the number of utterances (in both conditions, the Challenger speaks slightly more than the OP), or of the success of the conversation (no statistical difference in $\Delta$’s). This result complements the above-mentioned studies by providing insights into how a speaker’s situational role in a conversation mediates their influence on its dynamics.
@@ 7 Conclusion
In this work we introduce a similarity measure for conversational dynamics and develop a validation procedure to compare different representations and alignment methods. We showcase the measure’s applications in the context of an online debate community, adding to the literature on the relation between situational power and influence in conversations.
Our measure joins a growing toolkit of computational methods for conversational analysis. To encourage its application to other application domains, we make our code publicly available.
Finally, our measure could be used towards a more holistic evaluation of LLMs’ conversational ability, going beyond the quality of each individual reply to compare systems with respect to the dynamics they engender. Similarity measures like ConDynS can also be a step towards providing conversational-level feedback to AI agents to encourage dynamics that are similar to those preferred by humans (e.g., by extending RLHF).
@@ 8 Limitations
ConDynS should not be regarded as a conclusive measure for conversation dynamic similarity, but as a starting point for better approaches. It relies on simple prompting for multiple components, and each of them includes non-trivial tasks. Specifically, we noticed the difficulty of quantifying the alignment of two conversational dynamics. Without a specific rubric, it is difficult to interpret the score outputted by a model. While our measure provides a short description of the analysis for interpretability, there is much room for future work to systematize the scoring standard and procedure.
@@
More broadly, our analysis is limited to a single domain, which was particularly convenient for developing and validating the measure. Various components (such as the prompts) might need to be adapted in order to apply the method to vastly different domains. By releasing our code in a modular fashion, we encourage adaptations and applications to other domains.
In particular, we only applied ConDynS to online text-based conversations. Future work could study conversation similarity in multi-modal contexts, exploring how to capture dynamics carried out through audio (e.g., voice inflection, tone) or visual (e.g., expressions, gestures) cues and aligning those dynamics.
Finally, our results provide new insights into the role of situational power in conversations. While in our analysis we control for speaker-specific factors, such as demographics, future work could explore what characteristics beyond the role of the speaker in conversation mediate their influence on the dynamics. Furthermore, combining our measure with a controlled experiment could complement our observational study to elucidate the causal link between situational power and conversational dynamics.
@@ Acknowledgments
We are grateful for engaging discussions with Team Zissou—including Yash Chatha, Nicholas Chernogor, Tushaar Gangavarapu, Laerdon Yah-Sung Kim, Lillian Lee, Vivian Nguyen, Luke Tao, Son Tran, and Ethan Xia. This work was enabled by a Gemma Academic Program GCP Credit Award. Cristian Danescu-Niculescu-Mizil was funded in part by the U.S. National Science Foundation under Grant No. IIS-1750615 (CAREER) and by Cornell’s Center for Social Sciences. Any opinions, findings, and conclusions in this work are those of the author(s) and do not necessarily reflect the views of Cornell University or the National Science Foundation.
@@ Appendix A Prompts for ConDynS
In this section, we provide the prompts used in ConDynS. We arrived on these prompts after exploring multiple different versions of the prompt and qualitatively examining the results.
SCD generation. We used the procedural prompt presented in Hua et al. (2024). The prompt is presented in Figure 5.
@@
SoP generation. We prompt a LLM to parse a SCD into a sequence of patterns (SoP). The prompt is presented in Figure 6.
@@
The model is asked to provide a score and a short description of the analysis for each pattern in Python dictionary format. The prompt can be found at Figure 7.
@@ Appendix B Examples of SCD and SoP
Table 3 shows three examples of SCD and SoP representation. They are machine-generated SCD and SoP of real conversations from CMV dataset. The first two conversations are very similar; the last is very different from the first in terms of dynamics. The entire transcript of each conversation can be found in Figure 8, Figure 9, and Figure 10.
@@
Table 3: Example SCD and SoP representations of three conversations. Conversation 2 (colored in blue) is a similar conversation (positive) of Conversation 1 (ConDynS assign score of 0.544). Conversation 3 (colored in red) is a non-similar conversation (negative) of Conversation 1 (score of 0.112). See Figure 8, 9, and 10 for the entire transcript of each conversation.
@@ C.1 Baseline Implementation Details
Cosine similarity. We use a pre-trained sentence BERT model ‘all-MiniLM-L6-v2’ (22.722.722.7M parameter in size) to map either the entire transcript or the generated SCD of the conversation into a 384 dimensional dense vector space. Text that is longer than 256 tokens are truncated. The similarity between two conversations is measured by calculating the cosine similarity of their embeddings.
BERTScore. We use a distilled version of the BERT base model ( [^67] M parameters) Sanh et al. (2020) and Huggingface’s BERTScore pipeline to calculate the similarity score.
@@ C.3 Topic setting during simulation
For the same topic setting, the topic identified for an anchor conversation is used as the imposed topic to simulate its positive and negative counterparts. For the different topic setting, topics identified from the 505050 conversations in the dataset are first shuffled. Each anchor is then assigned one of these shuffled topics, which serves as the specified topic for simulating its positive and negative counterparts, while ensuring that the shuffle topic is different from the anchor’s original identified topic. For the adversarial setting, a topic obtained through the shuffling process described in the different topic setting is used for simulating the positive counterpart, while the anchor’s original identified topic is used for simulating the negative counterpart.
@@ Appendix E Qualitative Examples
Table 5 includes multiple examples of identified patterns in each cluster.
@@
shift to ligher tone SPK2 shifts to a more agreeable tone.
@@
Table 5: Qualitative analysis of the two identified clusters. Dynamics found in Cluster 1 are on the top; dynamics found in Cluster 2 are on the bottom. Phrases in parentheses are distinguishing words used during the analysis.
@@ Appendix F Additional Applications Results
Figure 16 demonstrates the intra-group similarity between set $\Delta$ and set $\neg\Delta$, as described in Section 2

### v2

@@ 6.2 New investigation: speakers’ tendencies
The dynamics of a conversation are the result of a joint process involving all speakers. However, similar to how speakers have a tendency to use a specific style across different conversations Welch et al. (2019); Zhang et al. (2020a), they may also have a tendency to engage in certain types of dynamics. In fact, we find evidence to that effect: conversations that share a common speaker have more similar dynamics than those which do not (ConDynS $0.37$ vs. $0.35$, $p<0.001$ according to Mann Whitney U-test, comparing similarities within two samples of 1,000 pairs each).
@@
Alternatively, prior studies emphasize the critical role of persuasion strategies in debates and their outcomes Braca and Dondio (2023); Orazi et al. (2025). The Challenger, by selecting these strategies, may dictate the dynamics.
To distinguish between these two hypotheses, we design a setup that controls for speaker-related confounds, such as demographics that might otherwise spuriously correlate with both their role in the conversation and their influence (Figure 4). We select speakers who participate in at least four conversations, each started by a different post: two in which they take the role of OP and two in which they are the Challenger. [^8] There are $482$ such speakers, considering only dyadic conversations that pass our strict length filters.
We find that the pair in which the speaker is the OP is more similar than the pair in which they are the Challenger ($0.41$ vs. $0.37,p<0.01$ according to Wilcoxon signed-rank test). This suggests that conversation dynamics are more likely to follow the tendencies of the (higher-powered) OP than those of the Challenger, supporting the first hypothesis. [^9] We note that this result is not a mere consequence of the number of utterances (in both conditions, the Challenger speaks slightly more than the OP), or of the success of the persuasion attempt (no statistical difference in $\Delta$’s). This result complements the above-mentioned studies by providing insights into how a speaker’s situational role in a conversation mediates their influence on its dynamics.
7 Application to Other Settings
In developing ConDynS, we purposely adopted a flexible framework that can accommodate a broad range of conversational dynamics (Section 4). This design choice facilitates the extension of our measure to new domains with diverse dynamics.
Casual dialogues between friends. To explore this versatility, we apply our method to 50 dialogues from the Friends TV show Chen and Choi (2016) which are scripted to resemble casual face-to-face dialogues. The only component of ConDynS that needs to be adapted is the prompt for generating the SCDs. To tailor these summaries to the dynamics of this particular setting, we adjust the prompt to indicate its casual nature, incorporating a few hand-written examples of summaries (Figure 7 in the Appendix). We apply the same procedure from Section 6 to provide a qualitative analysis of the two top-level clusters obtained using ConDynS in this setting. The clusters (Table 9 in the Appendix) capture a contrast between (1) more serious conversations, where participants disclose vulnerabilities to seek validation or provide reassurance, sometimes using humor defensively to deflect tension, and (2) more lighthearted interactions, where humor and inside jokes are employed to playfully tease one another.
Non-English conversations. To examine how our method extends beyond English, we analyze 100 discussions from the German Wikipedia Talk Pages Hua et al. (2018). In addition to example hand-written summaries from this dataset, we add to the SCD prompt an instruction to generate the summaries in English. The resulting clusters are intuitive (Table 10 in the Appendix), distinguishing between (1) conversations in which disagreements are resolved through polite guidance supported by evidence and rationales and (2) conversations where disagreements remain persistent and are marked by sarcasms and accusations.
Overall, this exploration suggests ConDynS is versatile enough to be adapted to a variety of conversational settings and to account for different types of conversational dynamics. The ability of ConDynS to take into account specific aspects of conversational dynamics ultimately depends on the extent to which the SCD-creation procedure can be guided towards surfacing those aspects.
@@ 8 Conclusion
In this work, we introduce a similarity measure for conversational dynamics and develop a validation procedure to compare different representations and alignment methods. We showcase the measure’s utility in the context of an online debate community, adding to the literature on the relation between situational power and influence in conversations. Our measure joins a growing toolkit of computational methods for conversational analysis.
In future work, our measure could be used towards a more holistic evaluation of LLMs’ conversational ability, going beyond the quality of each individual reply to compare systems with respect to the dynamics they engender. For instance, our metric could be applied to compare the dynamics of mental health therapy conversations held by human therapists versus AI therapists, a timely problem with important societal ramifications.
Similarity measures like ConDynS can also be an important step towards providing conversational-level feedback to AI agents to encourage dynamics that are similar to those preferred by humans. For example, post-training reinforcement-learning methods could be extended to consider rewards based on similarity with human-preferred dynamics, in addition to human-preferred replies.
@@ 9 Limitations
ConDynS should not be regarded as a conclusive measure for conversation dynamic similarity, but as a starting point for better approaches. It relies on simple prompting for multiple components, and each of them includes non-trivial tasks. Specifically, we noticed the difficulty of quantifying the alignment of interactional dynamics. Without a specific rubric, it is difficult to interpret the score that the model outputs. While our measure provides a short description of the analysis for interpretability, there is much room for future work to systematize the scoring standard and procedure.
@@
More broadly, our main analysis is focused on a single domain, which was particularly convenient for developing and validating the measure. As shown in our exploratory analysis of additional domains, the prompt for generating SCD needs to be adapted in order to apply the method to different settings and languages. The usefulness of the comparison provided by ConDynS hinges on the type of dynamics captures by the SCDs. Further work is needed to explore the limits of SCD generation in vastly different conversational settings, and to develop more sophisticated methods that can better capture relevant dynamics.
By releasing our code in a modular fashion, together with demos on several domains, we encourage adaptations and applications to other domains. A particular challenge is translating our method to multi-modal contexts, exploring how to compare dynamics emerging from vocal (e.g., voice inflection, tone) or visual (e.g., expressions, gestures) patterns.
Finally, our analysis provides new insights into the role of situational power in conversations. While in our analysis we control for speaker-specific factors, such as demographics, future work could explore what characteristics beyond the role of the speaker in conversation mediate their influence on the dynamics. Furthermore, combining our measure with a controlled experiment could complement our observational study to elucidate the causal link between situational power and conversational dynamics.
@@ Acknowledgments
We thank the reviewers for their feedback, which in one case resulted in a conversation with dynamics that are hard to match. We are grateful for engaging discussions with Team Zissou—including Yash Chatha, Nicholas Chernogor, Tushaar Gangavarapu, Laerdon Yah-Sung Kim, Lillian Lee, Vivian Nguyen, Luke Tao, Son Tran, and Ethan Xia. This work was enabled by a Gemma Academic Program GCP Credit Award. We gratefully acknowledge use of the research computing resources of the Empire AI Consortium, Inc, with support from Empire State Development of the State of New York, the Simons Foundation, and the Secunda Family Foundation. Cristian Danescu-Niculescu-Mizil was funded in part by the U.S. National Science Foundation under Grant No. IIS-1750615 (CAREER), by Cornell’s Center for Social Sciences, by a LinkedIn Research Award, and by a Wikimedia Research Fund Award. Any opinions, findings, and conclusions in this work are those of the author(s) and do not necessarily reflect the views of Cornell University or the National Science Foundation.
@@ Appendix A Prompts for ConDynS
In this section, we provide the prompts used in ConDynS for all three domains where we applied the measure. We arrived on these prompts after exploring multiple different versions of the prompt and qualitatively examining the results.
SCD generation. To generate SCDs for ChangeMyView subreddit conversations, we used the procedural prompt presented in Hua et al. (2024). The prompt is presented in Figure 5. For conversations from German Wikipedia talk-pages, we adapt the prompt to focus on collaboration, as presented in Figure 6. Friends dialogues focus on sentiment and individual characters’ intentions, thus the prompt is modified as shown in Figure 7.
@@
Write a short summary capturing the trajectory of a Wikipedia talk-page discussion. Do not include specific article content, titles, policy names, diffs/edits, quotes, or concrete claims. The style you should avoid: Example Sentence 1: “Speaker1 insisted an article include a particular detail and cited a specific policy by name. Speaker2 countered with a different policy and argued that the section should be removed. Speaker3 referenced a prior version and proposed a precise rewrite.” Instead, do include indicators of sentiments (e.g., sarcasm, politeness, frustration), intentions (e.g., agreement, disagreement, rebuttal, concession, clarification, accusation), and strategies (e.g., consensus attempts, moderation, revert-restore cycles, rhetorical questions, appeals to emotion). The following sentences demonstrate the style you should follow: Example Sentence 2: “Both speakers hold differing views and become defensive. Speaker1 diminishes the weight of Speaker2’s reasoning, and Speaker2 blames Speaker1 for an uncivil tone. Both accuse each other of focusing on personal traits rather than reasoning.” Example Sentence 3: “The speakers refute each other with back-and-forth accusations. Persistent fault-finding and critical stances escalate tension and hinder productive discussion.” Overall, the trajectory summary should capture the key moments where the discussion’s tone or coordination changes. Here is an example of a complete trajectory summary. Trajectory Summary:
Multiple speakers discuss possible changes. Several present differing stances in sequence, building on and contesting each other’s reasoning. Speaker1 disputes a point from Speaker2, prompting a rebuttal. Speaker3 supports Speaker1, after which Speaker2 defends their position. Later, a speaker references a removed remark and offers an extended counter. Despite friction, the tone remains mostly civil with attempts at consensus.
Now, provide the trajectory summary for the following conversation. Conversation Transcript:
Now, summarize this conversation. Remember, do not include specific topics, claims, policies, or edits. Instead, capture the speakers’ sentiments, intentions, and strategies. Limit the trajectory summary to 80 words.
Trajectory Summary (in English):
Figure 6: Procedural prompt for generating SCD on German Wikipedia talk-page discussions
Write a short summary capturing the trajectory of a casual conversation. Do not include specific topics, events, or arguments from the conversation. The style you should avoid: Example Sentence 1: “Speaker1 said they had a difficult day at work, and mentioned that their boss was unfair. Speaker2 listened and agreed that bosses can be tough, then suggested they go out for dinner to forget about it..” Instead, do include indicators of sentiments (e.g., warmth, empathy, humor, nostalgia, vulnerability, support), individual intentions (e.g., building rapport, offering reassurance, seeking validation, self-disclosure, active listening, gentle disagreement, creating distance), and conversational strategies (if any) such as “collaborative storytelling”, “inside jokes”, “mirroring emotions”, and “affectionate teasing”. The following sentences demonstrate the style you should follow: Example Sentence 2: “Both speakers have similar feelings and appeared mutually supportive. Speaker1 initiates with a moment of self-disclosure, and Speaker2 responds with empathy and validation. Both speakers build on this exchange, strengthening their rapport.” Example Sentence 3: “The two speakers connected with back-and-forth affectionate teasing. Throughout the conversation, they kept building on each other’s humor with playful remarks, creating a lighthearted and comfortable discussion.” Overall, the trajectory summary should capture the key moments where the emotional connection of the conversation notably changes. Here is an example of a complete trajectory summary. Trajectory Summary: The conversation begins with two speakers exchanging neutral, surface-level comments. Speaker1 then shifts the tone by sharing a personal anecdote, prompting Speaker2 to respond with warmth and empathy. Speaker1 elaborates on their story and their need, but Speaker2 does not extend their support but retracts it. Now, provide the trajectory summary for the following conversation. Conversation Transcript: Now, summarize this conversation. Remember, do not include specific topics, claims, or arguments from the conversation. Instead, try to capture the speakers’ sentiments, intentions, and conversational/persuasive strategies. Limit the trajectory summary to 80 words. Trajectory Summary:
Figure 7: Procedural prompt for generating SCD on Friends conversations
SoP generation. We prompt a LLM to parse a SCD into a sequence of patterns (SoP). The prompt is presented in Figure 8. With the prompt being highly generic, no adaptation is necessary across conversation domains.
@@
The model is asked to provide a score and a short description of the analysis for each pattern in Python dictionary format. The prompt can be found at Figure 9. The prompt can be used across conversation settings.
@@ Appendix B Examples of SCD and SoP
Table 3 shows three examples of SCD and SoP representation. They are machine-generated SCD and SoP of real conversations from CMV dataset. The first two conversations are very similar; the last is very different from the first in terms of dynamics. The entire transcript of each conversation can be found in Figure 10, 11, and 12.
@@
Table 3: Example SCD and SoP representations of three conversations. Conversation 2 (colored in blue) is a similar conversation (positive) of Conversation 1 (ConDynS assign score of 0.544). Conversation 3 (colored in red) is a non-similar conversation (negative) of Conversation 1 (score of 0.112). See Figure 10, 11, and 12 for the entire transcript of each conversation.
@@ C.1 Baseline Implementation Details
Cosine similarity. We use a pre-trained sentence BERT model ‘all-MiniLM-L6-v2’ (22.722.7M parameter in size) to map either the entire transcript or the generated SCD of the conversation into a 384 dimensional dense vector space. Text that is longer than 256 tokens are truncated. The similarity between two conversations is measured by calculating the cosine similarity of their embeddings.
BERTScore. We use a distilled version of the BERT base model (6767M parameters) Sanh et al. (2020) and Huggingface’s BERTScore pipeline to calculate the similarity score.
@@ C.3 Topic setting during simulation
For the same topic setting, the topic identified for an anchor conversation is used as the imposed topic to simulate its positive and negative counterparts. For the different topic setting, topics identified from the $50$ conversations in the dataset are first shuffled. Each anchor is then assigned one of these shuffled topics, which serves as the specified topic for simulating its positive and negative counterparts, while ensuring that the shuffle topic is different from the anchor’s original identified topic. For the adversarial setting, a topic obtained through the shuffling process described in the different topic setting is used for simulating the positive counterpart, while the anchor’s original identified topic is used for simulating the negative counterpart.
@@ Appendix E Qualitative Examples
Table 5 and Table 6 includes multiple examples of identified patterns in each cluster from the CMV conversations selected with strict length control as reported in the main paper. Table 7 and Table 8 includes examples of patterns from each cluster of CMV conversations selected with less strict length control, as discussed in Appendix F. Table 9 includes examples and identified patterns in each cluster from the Friends dialogues and Table 10 includes those from the German Wikipedia talk-pages conversations.
@@
shift to lighter tone SPK2 shifts to a more agreeable tone.
@@
Table 5: Qualitative analysis of dynamics of Cluster 1 from ChangeMyView conversations (with strict length control). Phrases in parentheses are distinguishing words used during the analysis.
# Category Dynamics Examples
@@
Table 6: Qualitative analysis of dynamics of Cluster 2 from ChangeMyView conversations (with strict length control). Phrases in parentheses are distinguishing words used during the analysis.
@@ Appendix F Additional Applications Results
Application result without strict length control. We report here the experiment results equivalent to those presented in Section 6 without the strict length control (maintaining all the other filters).
We conduct clustering and then compare inter-group similarity as well as intra-group diversity (outlined in Figure 18).
Figure 18: Outline for applying ConDynS to different analyses supported by similarity measures with less strict length control. Statistical significant differences marked with ∗∗∗ ($p<0.001$).
We first cluster a random sample of 200 conversations from the year 2018, using hierarchical clustering with ConDynS. The qualitative analysis of the two top-level clusters, focusing on aggregated patterns with alignment scores $s_{i}>0.5$, reveals a similar separation in the use of tone, conversational strategies, and evolving dynamics compared to the results obtained under stricter length control. Illustrative examples of the corresponding patterns are provided in Table 7 and Table 8. We also observe distinctions between the two top-level clusters when considering persuasion success labels ($\Delta$). In this random set, 13% of conversations are labeled with $\Delta$. Cluster 1 and Cluster 2 differ significantly in proportion of conversations receiving $\Delta$ (35% vs. 3%, $p<0.0001$ according to z-test for proportions). We further sample 100 conversations labeled with $\Delta$ (set $\Delta$) and 100 without (set $\neg\Delta$). We find that conversations in Cluster 1 are more similar to set $\Delta$ than to set $\neg\Delta$ (mean ConDynS 0.43 vs. 0.34, $p<0.001$, Mann–Whitney U-test). Moreover, set $\Delta$ conversations are significantly more similar to each other than set $\neg\Delta$ conversations (mean ConDynS 0.53 vs. 0.37, $p<0.001$, Mann–Whitney U-test).
To investigate speakers’ tendencies, we compare the similarity between 1,000 pairs of conversations sharing a common speaker with the similarity between 1,000 pairs without a shared speaker, finding that that pairs sharing a common speaker demonstrate more similar dynamics than those without a shared speaker (ConDynS 0.39 vs. 0.37, $p<0.01$, Mann–Whitney U-test). We then focus on the 486 speakers who each participated in at least four conversations—two as the OP and two as the Challenger—and we further observe that pairs where the speaker serves as OP are more similar than pairs where the same speaker serves as Challenger (0.41 vs. 0.38, $p<0.05$ according to a Wilcoxon signed-rank test). The setup described in Figure 19.
Figure 19: Similarity between two conversations in which a speaker has the role of OP vs. between two conversations in which the same speaker is the Challenger. The difference is statistically significant ($p<0.05$). Conversations are filtered with less strict length control.
This again suggests that conversational dynamics are more strongly shaped by the OP than by the Challenger, aligning with our findings in the main paper.
As shown in this section, we found the results to be both qualitatively and quantitatively similar to those reported in the main paper under stricter conversation length control (restricting to conversations of 4–6 utterances).
# Category Dynamics Examples
1 Tone negative politeness SPK1 then concedes, acknowledging a shift in understanding and expressing gratitude.
(gratitude, thanks, appreciation) SPK2 concludes with gratitude and agreement.
collaborative The conversation concludes with mutual respect and acknowledgment, with both speakers reflecting on their own viewpoints.
(collaborative, shared) SPK2 concludes by asserting a shared understanding, attempting to resolve the perceived disagreement and establish common ground.
conciliatory SPK1 concedes that SPK2’s explanation has shifted their perspective.
(acknowledges, concede) SPK1 acknowledges their error.
Strategy elaboration SPK2 elaborates on their reasoning and acknowledges SPK1’s potential correctness, demonstrating a degree of concession.
(elaborate, detailed) SPK2 offers their own reasons for disliking the episode, providing a detailed explanation.
agreement SPK1 expresses agreement.
(agreement) SPK2 concedes to a point made by SPK1, expressing agreement.
compromise SPK1 responds by agreeing with one aspect of SPK2’s statement while also introducing a contrasting viewpoint.
(compromise) SPK1 concedes a point but expresses a pessimistic view of consumer behavior.
Changes changes in perspective SPK1 expresses gratitude for the information, indicating a change in their understanding.
(revised, change) SPK1 acknowledges the potential negative consequences and concedes, changing their view.
shift to ligher tone The tone shifts from inquisitive to reflective and ultimately appreciative.
The tone shifts to a friendly and helpful exchange.
Table 7: Qualitative analysis of Cluster 1 dynamics from ChangeMyView conversations (selected with less strict length control). Phrases in parentheses are distinguishing words used during the analysis.
# Category Dynamics Examples
2 Tone dismissive SPK2 initiates the conversation with an accusatory and dismissive tone, directly attacking SPK1’s reasoning.
(accusatory, dismissive) SPK2 dismisses the example as irrelevant to their point.
sarcastic SPK1 refutes SPK2’s claims, employing sarcasm.
(sarcasm, sarcastic) SPK1 expresses a sense of resignation, possibly sarcastic.
defensive The tone shifts from informative to defensive.
(defensive, refute) SPK1 responds defensively, attempting to clarify their position and refute SPK2’s interpretation.
confrontational SPK1 accuses SPK2 of sexism and attributes historical disparities to societal constraints.
(accuses, blame) SPK2 starts the conversation with a rhetorical question, implying blame.
Strategy straw man fallacy SPK1 accuses SPK2 of using a straw man fallacy.
(straw man) SPK2 suggests a limited perspective and employs a straw man fallacy.
providing examples SPK2 provides examples to support their argument.
(examples) SPK1 rebuts SPK2’s points by dismissing anecdotal evidence.
analogy SPK2 then uses an analogy to challenge SPK1 reasoning.
(analogy) SPK2 then uses an analogy to further clarify their position.
seeking clarification SPK1 responds defensively, seeking clarification.
(confusion, seeking clarification) SPK2 expresses confusion and presses SPK1 to define the specific group to which this obligation applies.
disagreement SPK2 continues to disagree and questions SPK1’s perspective.
(disagrees, contrasting) SPK1 immediately expresses disagreement, employing a comparison to other controversial industries to undermine SPK2 claims.
direct responses SPK1 directly answers the question with disagreement, citing practical concerns.
(direct, immediately)
Changes maintains perspective SPK1 denies the accusation and reiterates their stance.
(reiterates, persists in) SPK1 persists in their disagreement, providing counter-evidence.
shift to contentious tone The tone shifts from informative to defensive.
The tone shifts from neutral inquiry to a more challenging and potentially critical stance.
Table 8: Qualitative analysis of Cluster 2 dynamics from ChangeMyView conversations (selected with less strict length control). Phrases in parentheses are distinguishing words used during the analysis.
@@
Additional result. Figure 20 demonstrates the intra-group similarity between set $\Delta$ and set $\neg\Delta$, as described in Section 2.